OpenStudy (anonymous):

solve the equation by combining like terms x/2+x^2/3-x/5-x^2/4

OpenStudy (anonymous):

solve the equation by combining like terms x/2+x^2/3-x/5-x^2/4

OpenStudy (saifoo.khan):

\[\frac x2+\frac{x^2}{3}-\frac x5-\frac{x^2}4\]Like that?

OpenStudy (anonymous):

yea

OpenStudy (saifoo.khan):

Do you know how to find LCD?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ive never done these types of problems before

OpenStudy (anonymous):

thanxs

OpenStudy (anonymous):

do i have to find all the least common multiples for all four numbers

OpenStudy (saifoo.khan):

Yes! (:

OpenStudy (anonymous):

well the least common multiple of 2 and 3 and 4 is 8 but five doesnt have a common multiple of 8

OpenStudy (anonymous):

can you help me im still lost

OpenStudy (saifoo.khan):

2*2*3*5